Will you please help me with the homework below.
consider the statement below and write each step as as algebraic expression
Step1: Think of a number
Step2: Add 7
Step3: Double your answer
Step4: Subtract 4
Step5: Divide by 2
Step6: Subtract the number you started
Step7: write down your answer
Thank you

Found 2 solutions by ewatrrr, greenestamps:
Answer by ewatrrr(24785)


Hi
Think of a number 'n'
 (2(n+7) - 4)/2 - n =  (2n + 10)/2 - n = n + 5 - n = 5
Fun anyway!
Wish You the Best in your Studies.


Answer by greenestamps(13200)


(A) Without simplifying along the way....
Step1: Think of a number ---------------- x Step2: Add 7 ---------------------------- x+7 Step3: Double your answer --------------- 2(x+7) Step4: Subtract 4 ----------------------- 2(x+7)-4 Step5: Divide by 2 ---------------------- (2(x+7)-4)/2 Step6: Subtract the number you started -- ((2(x+7)-4)/2)-x Step7: write down your answer
(B) Simplifying as we go....
Step1: Think of a number ---------------- x Step2: Add 7 ---------------------------- x+7 Step3: Double your answer --------------- 2(x+7) = 2x+14 Step4: Subtract 4 ----------------------- (2x+14)-4 = 2x+10 Step5: Divide by 2 ---------------------- (2x+10)/2 = x+5 Step6: Subtract the number you started -- (x+5)-x = 5 Step7: write down your answer ----------- 5